OpenStudy (anonymous):

Consider the solid obtained by rotating the region bounded by the given curves about the y-axis. y = ln(x), y = 1, y = 5, x = 0 Find the volume V of this solid.

OpenStudy (dumbcow):

ok this time the cross-sections are horizontal because we are rotating about the y-axis Since we are rotating the shape of each cross-section is a circle Area of circle is: \[\pi*r^{2}\] the radius can be represented by the x-distance the line ln(x) is away from y-axis \[y= \ln x \rightarrow x = e^{y}\] so \[r =e^{y}\] Integral looks like this \[V = \pi \int\limits_{1}^{5}(e^{y})^{2} dy\]